The demand for calculators can be represented by PD = 120 - (1/4)Q, and the supply of oranges is represented by pS = 30 + (1/The demand for calculators can be represented by PD = 120 - (1/4)9, and the supply of oranges is represented by pS = 30 + (1/The demand for calculators can be represented by PD = 120 – (1/4)Q, and the supply of oranges is represented by pS = 30 + (1/The demand for calculators can be represented by PD = 120 - (1/4)9D, and the supply of oranges is represented by pS = 30 + (1The demand for calculators can be represented by PD = 120 - (1/4)QD, and the supply of oranges is represented by pS = 30 + (1
Answer #1

Answer 1. $1800

reason- Consumer surplus=0.5(120-90)(120)= $1800

Answer 2. 48

reason- World price=$80

tariff=$2. new world price= $82

82=120-0.25Qd

Qd=152

82=30+0.5Qs

Qs=104

Import= 152-104= 48

Answer 3. $3600

reason- Producer surplus=0.5(90-30)(120)=$3600

Answer 4. Equilibrium Price=$90

Answer 5. Equilibrium quantity=120

Reason-

Equilibrium

120-0.25Q=30+0.5Q

Equilibrium quantity Q*= 120

Equilibrium Price P*= 30+0.5*120=$90